Shift Summary for owl shift, 01/05/2000


Subject: Shift Summary for owl shift, 01/05/2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Wed Jan 12 2000 - 16:39:31 CST


              Shift Summary for owl shift, 01/05/2000

On Shift: J. Shields, Bob T.

Hours of beam available at all intensities: 0.0
Hours of beam available at nominal intensity: 0.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 4.5 (cosmics)

Accounting of lost time:

  magnet replacement, Tevatron RF and P1 line power supply problems

========================================================================
Runs ended this shift:

Run 15398 COSMICS

Tapes brought to FCC this shift:
  (including tapes for online split)

none

========================================================================
Accesses during shift:

one. (Bob T. made measurements of the beam holes)

========================================================================
Overview of shift:
 
ran cosmics, waited impatiently for beam, and went on a veeerry brief access

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15398 00:00 beam is down for replacement of a magnet.
                         Estimate is 03:00.

              
              03:21 new estimate is 04:30

15399 04:20 stop cosmics run and initialize a muon run
                        in order to be ready for beam

              04:38 Ch 13 says, "Beginning Main Injector turn-on.
                       Estimate for beam 0500"

              04:45 RT, JS go downstairs on access. Bob T. uses a
                      tape measure to make ultra-precise measurements
                      of the beam holes.

              04:47 JS, RT return from the strenuous and grueling access.

              05:01 no beam. P1 line power supply problems.
                       estimate 5:30

              05:34 new estimate is 06:05

              06:03 Now they're investigating TeV RF problems.
                       new estimate is 06:30

              06:30 no beam. No update from Chan 13 yet.
      
              06:50 shift crew begins to struggle to come to terms
                      with the harsh reality that they are unlikely
                      to get beam on their watch.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift



This archive was generated by hypermail 2b27 : Wed Jan 12 2000 - 16:39:34 CST